if the sum of the series 1+ 2/x + 4/x^2 + 8/x^3 +...... infinity is a finite number then
(a) x
(b) x>1/2
(c) x>-2
(d) x2

Grade:12th pass

2 Answers

Y RAJYALAKSHMI
9 years ago
1+ 2/x + 4/x^2 + 8/x^3 +......  is sum of terms in GP where r = 2/x.  
The sum is finite if |r| |2/x| |x| > 2 => x 2
 
Ans: x 2
SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
The given series can be written as
1+ 2/x + 4/x^2 + 8/x^3 +......=(1-(2/x))-1
where r=2/x
For a series to be finite |r|<1 so |2/x|<1
So |x|>2